OpenStudy (anonymous):

4 2/3 - 3 3/5.. can u please show me the steps

OpenStudy (anonymous):

16/15

OpenStudy (anonymous):

\[4\frac{2}{3} - 3\frac{3}{5} = \frac{14}{3} - \frac{18}{5} = \frac{70 - 54}{15} = \frac{16}{15} = 1\frac{1}{15}\]